WritingTools-TheSTYLEandDICTIONPrograms L.L.Cherry W.Vesterman LivingstonCollege RutgersUniversity readability ABSTRACT Textprocessingsystemsarenowinheavyuseinmanycompaniestoformat documents.Withmanydocumentsstoredonline,ithasbecomepossibletouse computerstostudywritingstyleitselfandtohelpwritersproducebetterwritten andmorereadableprose.Thesystemofprogramsdescribedhereisaninitial steptowardsuchhelp.Itincludesprogramsandadatabasedesignedtoproduce astylisticprofileofwritingatthewordandsentencelevel.Thesystemmeasures readability,sentenceandwordlength,sentencetype,wordusage,andsentence openers.Italsolocatescommonexamplesofwordyphrasingandbaddiction. Thesystemisusefulforevaluatingadocument’sstyle,locatingsentencesthat maybedifficulttoreadorexcessivelywordy,anddeterminingaparticular writer’sstyleoverseveraldocuments. 8April1993 WritingTools-TheSTYLEandDICTIONPrograms L.L.Cherry W.Vesterman LivingstonCollege RutgersUniversity readability 1.Introduction Computershavebecomeimportantinthedocumentpreparationprocess,withprogramsto checkforspellingerrorsandtoformatdocuments.Astheamountoftextstoredonlineincreases, itbecomesfeasibleandattractivetostudywritingstyleandtoattempttohelpthewriterinproducingreadabledocuments.Thesystemofwritingtoolsdescribedhereisafirststeptowardsuch help.Thesystemincludesprogramsandadatabasetoanalyzewritingstyleatthewordandsentencelevel.Weusetheterm‘‘style’’inthispapertodescribetheresultsofawriter’sparticular choicesamongindividualwordsandsentenceforms.Althoughmanyjudgementsofstylearesubjective,particularlythoseofwordchoice,therearesomeobjectivemeasuresthatexpertsagree leadtogoodstyle.Threeprogramshavebeenwrittentomeasuresomeoftheobjectivelydefinablecharacteristicsofwritingstyleandtoidentifysomecommonlymisusedorunnecessary phrases.Althoughadocumentthatconformstothestylisticrulesisnotguaranteedtobecoherentandreadable,onethatviolatesalloftherulesislikelytobedifficultortedioustoread.The programSTYLEcalculatesreadability,sentencelengthvariability,sentencetype,wordusageand sentenceopenersatarateofabout400wordspersecondonaPDP11/70runningtheUNIX◯r OperatingSystem.Itassumesthatthesentencesarewell-formed,i.e.thateachsentencehasa verbandthatthesubjectandverbagreeinnumber.DICTIONidentifiesphrasesthatareeither badusageorunnecessarilywordy.EXPLAINactsasathesaurusforthephrasesfoundbyDIC-TION.Sections2,3,and4describetheprograms;Section5givestheresultsonacross-sectionof technicaldocuments;Section6discussesaccuracyandproblems;Section7givesimplementation details. 2.STYLE TheprogramSTYLEreadsadocumentandprintsasummaryofreadabilityindices,sentencelengthandtype,wordusage,andsentenceopeners.Itmayalsobeusedtolocateallsentencesinadocumentlongerthanagivenlength,ofreadabilityindexhigherthanagivennumber, thosecontainingapassiveverb,orthosebeginningwithanexpletive.STYLEisbasedonthesystemforfindingEnglishwordclassesorpartsofspeech,PARTS[1].PARTSisasetofprograms thatusesasmalldictionary(about350words)andsuffixrulestopartiallyassignwordclassesto Englishtext.Itthenusesexperimentallyderivedrulesofwordordertoassignwordclassestoall wordsinthetextwithanaccuracyofabout95%.BecausePARTSusesonlyasmalldictionary andgeneralrules,itworksontextaboutanysubject,fromphysicstopsychology.Stylemeasures havebeenbuiltintotheoutputphaseoftheprogramsthatmakeupPARTS.SomeofthemeasuresaresimplecountersofthewordclassesfoundbyPARTS;manyaremorecomplicated.For example,theverbcountisthetotalnumberofverbphrases.Thisincludesphraseslike:

-2-hasbeengoing wasonlygoing togo eachofwhicheachcountsasoneverb.Figure1showstheoutputofSTYLErunonapaperby KernighanandMasheyabouttheUNIXprogrammingenvironment[2]. programmingenvironment readabilitygrades: (Kincaid)12.3(auto)12.8(Coleman-Liau)11.8(Flesch)13.5(46.3) sentenceinfo: no.sent335no.wds7419 avsentleng22.1avwordleng4.91 no.questions0no.imperatives0 no.nonfuncwds436258.8%avleng6.38 shortsent(<17)35%(118)longsent(>32)16%(55) longestsent82wdsatsent174;shortestsent1wdsatsent117 sentencetypes: simple34%(114)complex32%(108) compound12%(41)compound-complex21%(72) wordusage: verbtypesas%oftotalverbs tobe45%(373)aux16%(133)inf14%(114) passivesas%ofnon-infverbs20%(144) typesas%oftotal prep10.8%(804)conj3.5%(262)adv4.8%(354) noun26.7%(1983)adj18.7%(1388)pron5.3%(393) nominalizations2%(155) sentencebeginnings: subjectopener:noun(63)pron(43)pos(0)adj(58)art(62)tot67% prep12%(39)adv9%(31) verb0%(1)subconj6%(20)conj1%(5) expletives4%(13) Figure1 Astheexampleshows,STYLEoutputisinfiveparts.Afterabriefdiscussionofsentences,we willdescribethepartsinorder. 2.1.Whatisasentence? Readersofdocumentshavelittletroubledecidingwherethesentencesend.Peopledon’t evenhavetostopandthinkaboutusesofthecharacter‘‘.’’inconstructionslike1.25,A.J.Jones, Ph.D.,i.e.,oretc..Whenacomputerreadsadocument,findingtheendofsentencesisnotas easy.Firstwemustthrowawaytheprinter’smarksandformattingcommandsthatlitterthetext incomputerform.ThenSTYLEdefinesasentenceasastringofwordsendinginoneof: .!?/. Theendmarker‘‘/.’’maybeusedtoindicateanimperativesentence.Imperativesentencesthat arenotsomarkedarenotidentifiedasimperative.STYLEproperlyhandlesnumberswith embeddeddecimalpointsandcommas,stringsoflettersandnumberswithembeddeddecimal pointsusedfornamingcomputerfilenames,andthecommonabbreviationslistedinAppendix1. Numbersthatendsentences,liketheprecedingsentence,causeasentencebreakifthenextword beginswithacapitalletter.Initialsonlycauseasentencebreakifthenextwordbeginswitha -3-capitalandisfoundinthedictionaryoffunctionwordsusedbyPARTS.Sothestring J.D.JONES doesnotcauseabreak,butthestring ...systemH.The... does.Withtheserulesmostsentencesarebrokenattheproperplace,althoughoccasionally eithertwosentencesarecalledoneorafragmentiscalledasentence.Moreonthislater. 2.2.ReadabilityGrades ThefirstsectionofSTYLEoutputconsistsoffourreadabilityindices.AsKlarepointsout in[3]readabilityindicesmaybeusedtoestimatethereadingskillsneededbythereadertounderstandadocument.ThereadabilityindicesreportedbySTYLEarebasedonmeasuresofsentence andwordlengths.Althoughtheindicesmaynotmeasurewhetherthedocumentiscoherentand wellorganized,experiencehasshownthathighindicesseemtobeindicatorsofstylisticdifficulty. Documentswithshortsentencesandshortwordshavelowscores;thosewithlongsentencesand manypolysyllabicwordshavehighscores.The4formulaereportedareKincaidFormula[4], AutomatedReadabilityIndex[5],Coleman-LiauFormula[6]andanormalizedversionofFlesch ReadingEaseScore[7].Theformulaedifferbecausetheywereexperimentallyderivedusingdifferenttextsandsubjectgroups.Wewilldiscusseachoftheformulaebriefly;foramoredetailed discussionthereadershouldsee[3]. TheKincaidFormula,givenby: ReadingGrade=11.8*sylperwd+.39*wdspersent−15.59 wasbasedonNavytrainingmanualsthatrangedindifficultyfrom5.5to16.3inreadinggrade level.Thescorereportedbythisformulatendstobeinthemid-rangeofthe4scores.Becauseit isbasedonadulttrainingmanualsratherthanschoolbooktext,thisformulaisprobablythebest onetoapplytotechnicaldocuments. TheAutomatedReadabilityIndex(ARI),basedontextfromgrades0to7,wasderivedto beeasytoautomate.Theformulais: ReadingGrade=4.71*letperwd+.5*wdspersent−21.43 ARItendstoproducescoresthatarehigherthanKincaidandColeman-Liaubutareusually slightlylowerthanFlesch. TheColeman-LiauFormula,basedontextrangingindifficultyfrom.4to16.3,is: ReadingGrade=5.89*letperwd−.3*sentper100wds−15.8 Ofthefourformulaethisoneusuallygivesthelowestgradewhenappliedtotechnicaldocuments. Thelastformula,theFleschReadingEaseScore,isbasedongradeschooltextcovering grades3to12.Theformula,givenby: ReadingScore=206.835−84.6*sylperwd−1.015*wdspersent isusuallyreportedintherange0(verydifficult)to100(veryeasy).Thescorereportedby STYLEisscaledtobecomparabletotheotherformulas,exceptthatthemaximumgradelevel reportedissetto17.TheFleschscoreisusuallythehighestofthe4scoresontechnicaldocuments. Coke[8]foundthattheKincaidFormulaisprobablythebestpredictorfortechnicaldocuments;bothARIandFleschtendtooverestimatethedifficulty;Coleman-Liautendtounderesti-mate.Ontextintherangeofgrades7to9thefourformulastendtobeaboutthesame.On easytexttheColeman-Liauformulaisprobablypreferredsinceitisreasonablyaccurateatthe lowergradesanditissafertopresenttextthatisalittletooeasythanalittletoohard.

-4-Ifadocumenthasparticularlydifficulttechnicalcontent,especiallyifitincludesalotof mathematics,itisprobablybesttomakethetextveryeasytoread,i.e.alowerreadabilityindex byshorteningthesentencesandwords.Thiswillallowthereadertoconcentrateonthetechnical contentandnotthelongsentences.Theusershouldrememberthattheseindicesareestimators; theyshouldnotbetakenasabsolutenumbers.STYLEcalledwith‘‘−rnumber’’willprintall sentenceswithanAutomatedReadabilityIndexequaltoorgreaterthan‘‘number’’. 2.3.Sentencelengthandstructure ThenexttwosectionsofSTYLEoutputdealwithsentencelengthandstructure.Almostall booksonwritingstyleoreffectivewritingemphasizetheimportanceofvarietyinsentencelength andstructureforgoodwriting.Ewing’sfirstruleindiscussingstyleinthebookWritingfor Results[9]is: ‘‘Varythesentencestructureandlengthofyoursentences.’’ Leggett,MeadandCharvatbreakthisruleinto3inPrentice-HallHandbookforWriters[10]as follows: ‘‘34a.Avoidtheoveruseofshortsimplesentences.’’ ‘‘34b.Avoidtheoveruseoflongcompoundsentences.’’ ‘‘34c.Usevarioussentencestructurestoavoidmonotonyandincreaseeffectiveness.’’ Althoughexpertsagreethattheserulesareimportant,notallwritersfollowthem.Sampletechnicaldocumentshavebeenfoundwithalmostnosentencelengthortypevariability.Onedocumenthad90%ofitssentencesaboutthesamelengthastheaverage;anotherwasmadeupalmost entirelyofsimplesentences(80%). Theoutputsectionslabeled‘‘sentenceinfo’’and‘‘sentencetypes’’givebothlengthandstructuremeasures.STYLEreportsonthenumberandaveragelengthofbothsentencesandwords, andnumberofquestionsandimperativesentences(thoseendingin‘‘/.’’).Themeasuresofnonfunctionwordsareanattempttolookatthecontentwordsinthedocument.InEnglishnonfunctionwordsarenouns,adjectives,adverbs,andnon-auxiliaryverbs;functionwordsarepreposi-tions,conjunctions,articles,andauxiliaryverbs.Sincemostfunctionwordsareshort,theytend tolowertheaveragewordlength.Theaveragelengthofnon-functionwordsmaybeamoreuseful measureforcomparingwordchoiceofdifferentwritersthanthetotalaveragewordlength.The percentagesofshortandlongsentencesmeasuresentencelengthvariability.Shortsentencesare thoseatleast5wordslessthantheaverage;longsentencesarethoseatleast10wordslongerthan theaverage.Lastinthesentenceinformationsectionisthelengthandlocationofthelongestand shortestsentences.Iftheflag‘‘−lnumber’’isused,STYLEwillprintallsentenceslongerthan ‘‘number’’. Becauseofthedifficultiesindealingwiththemanyusesofcommasandconjunctionsin English,sentencetypedefinitionsvaryslightlyfromthoseofstandardtextbooks,butstillmeasure thesameconstructionalactivity. 1.Asimplesentencehasoneverbandnodependentclause. 2.Acomplexsentencehasoneindependentclauseandonedependentclause,eachwithone verb.Complexsentencesarefoundbyidentifyingsentencesthatcontaineitherasubordinateconjunctionoraclausebeginningwithwordslike‘‘that’’or‘‘who’’.Theprecedingsentencehassuchaclause. 3.Acompoundsentencehasmorethanoneverbandnodependentclause.Sentencesjoined by‘‘;’’arealsocountedascompound. 4.Acompound-complexsentencehaseitherseveraldependentclausesoronedependentclause andacompoundverbineitherthedependentorindependentclause. Evenusingthesebroaderdefinitions,simplesentencesdominatemanyofthetechnicaldocumentsthathavebeentested,buttheexampleinFigure1showsvarietyinbothsentencestructure andsentencelength.

-5-2.4.WordUsage Thewordusagemeasuresareanattempttoidentifysomeotherconstructionalfeaturesof writingstyle.TherearemanydifferentwaysinEnglishtosaythesamething.Theconstructions differfromoneanotherintheformofthewordsused.Thefollowingsentencesallconveyapproximatelythesamemeaningbutdifferinwordusage: Thecxioprogramisusedtoperformallcommunicationbetweenthesystems. Thecxioprogramperformsallcommunicationsbetweenthesystems. Thecxioprogramisusedtocommunicatebetweenthesystems. Thecxioprogramcommunicatesbetweenthesystems. Allcommunicationbetweenthesystemsisperformedbythecxioprogram. Thedistributionofthepartsofspeechandverbconstructionshelpsidentifyoveruseofparticular constructions.AlthoughthemeasuresusedbySTYLEarecrude,theydopointoutproblem areas.Foreachcategory,STYLEreportsapercentageandarawcount.Inadditiontolookingat thepercentage,theusermayfinditusefultocomparetherawcountwiththenumberofsentences.If,forexample,thenumberofinfinitivesisalmostequaltothenumberofsentences,then manyofthesentencesinthedocumentareconstructedlikethefirstandthirdinthepreceding example.Theusermaywanttotransformsomeofthesesentencesintoanotherform.Someof theimplicationsofthewordusagemeasuresarediscussedbelow. Verbs aremeasuredinseveraldifferentwaystotrytodeterminewhattypesofverbconstructions aremostfrequentinthedocument.Technicalwritingtendstocontainmanypassiveverb constructionsandotherusageoftheverb‘‘tobe’’.Thecategoryofverbslabeled‘‘tobe’’ measuresbothpassivesandsentencesoftheform: subjecttobepredicate Incountingverbs,wholeverbphrasesarecountedasoneverb.Verbphrasescontaining auxiliaryverbsarecountedinthecategory‘‘aux’’.Theverbphrasescountedherearethose whosetenseisnotsimplepresentorsimplepast.Itmighteventuallybeusefultodomore detailedmeasuresofverbtenseormood.Infinitivesarelistedas‘‘inf’’.Thepercentages reportedforthesethreecategoriesarebasedonthetotalnumberofverbphrasesfound. Thesecategoriesarenotmutuallyexclusive;theycannotbeadded,since,forexample,‘‘to begoing’’countsasboth‘‘tobe’’and‘‘inf’’.Useofthesethreetypesofverbconstructions variessignificantlyamongauthors. STYLEreportspassiveverbsasapercentageofthefiniteverbsinthedocument.Most stylebookswarnagainsttheoveruseofpassiveverbs.Coleman[11]hasshownthatsentenceswithactiveverbsareeasiertolearnthanthosewithpassiveverbs.Althoughthe invertedobject-subjectorderofthepassivevoiceseemstoemphasizetheobject,Coleman’s experimentsshowedthatthereislittledifferenceinretentionbywordposition.Healso showedthatthedirectobjectofanactiveverbisretainedbetterthanthesubjectofapassiveverb.Theseexperimentssupporttheadviceofthestylebookssuggestingthatwriters shouldtrytouseactiveverbswhereverpossible.Theflag‘‘−p’’causesSTYLEtoprintall sentencescontainingpassiveverbs. Pronouns addcohesivenessandconnectivitytoadocumentbyprovidingback-reference.Theyare oftenashort-handnotationforsomethingpreviouslymentioned,andthereforeconnectthe sentencecontainingthepronounwiththewordtowhichthepronounrefers.Althoughthere areothermechanismsforsuchconnections,documentswithnopronounstendtobewordy andtohavelittleconnectivity. Adverbs canprovidetransitionbetweensentencesandorderintimeandspace.Inperformingthese -6-functions,adverbs,likepronouns,provideconnectivityandcohesiveness. Conjunctions provideparallelisminadocumentbyconnectingtwoormoreequalunits.Theseunitsmay bewholesentences,verbphrases,nouns,adjectives,orprepositionalphrases.Thecompoundandcompound-complexsentencesreportedundersentencetypeareparallelstruc-tures.Otherusesofparallelstructuresareindicatedbythedegreethatthenumberofcon-junctionsreportedunderwordusageexceedsthecompoundsentencemeasures. NounsandAdjectives. Aratioofnounstoadjectivesnearunitymayindicatetheover-useofmodifiers.Sometech-nicalwritersqualifyeverynounwithoneormoreadjectives.Qualifiersinphraseslike‘‘sim-plelinearsingle-linknetworkmodel’’oftenlendmoreobscuritythanprecisiontoatext. Nominalizations areverbsthatarechangedtonounsbyaddingoneofthesuffixes‘‘ment’’,‘‘ance’’,‘‘ence’’,or ‘‘ion’’.Examplesareaccomplishment,admittance,adherence,andabbreviation.Whena writertransformsanominalizedsentencetoanon-nominalizedsentence,she/heincreases theeffectivenessofthesentenceinseveralways.Thenounbecomesanactiveverbandfrequentlyonecomplicatedclausebecomestwoshorterclauses.Forexample, Theirinclusionofthisprovisionisadmissionoftheimportanceofthesystem. Whentheyincludedthisprovision,theyadmittedtheimportanceofthesystem. Colemanfoundthatthetransformedsentenceswereeasiertolearn,evenwhenthetransformationproducedsentencesthatwereslightlylonger,providedthetransformationbrokeone clauseintotwo.Writerswhofindtheirdocumentcontainsmanynominalizationsmaywant totransformsomeofthesentencestouseactiveverbs. 2.5.Sentenceopeners Anotheragreeduponprincipleofstyleisvarietyinsentenceopeners.BecauseSTYLE determinesthetypeofsentenceopenerbylookingatthepartofspeechofthefirstwordinthe sentence,thesentencescountedundertheheading‘‘subjectopener’’maynotallreallybeginwith thesubject.However,alargepercentageofsentencesinthiscategorystillindicateslackofvarietyinsentenceopeners.Othersentenceopenermeasureshelptheuserdetermineifthereare transitionsbetweensentencesandwherethesubordinationoccurs.Adverbsandconjunctionsat thebeginningofsentencesaremechanismsfortransitionbetweensentences.Apronounatthe beginningshowsalinktosomethingpreviouslymentionedandindicatesconnectivity. Thelocationofsubordinationcanbedeterminedbycomparingthenumberofsentencesthat beginwithasubordinatorwiththenumberofsentenceswithcomplexclauses.Iffewsentences startwithsubordinateconjunctionsthenthesubordinationisembeddedorattheendofthecomplexsentences.Forvarietythewritermaywanttotransformsomesentencestohaveleadingsubordination. Thelastcategoryofopeners,expletives,iscommonlyoverworkedintechnicalwriting. Expletivesarethewords‘‘it’’and‘‘there’’,usuallywiththeverb‘‘tobe’’,inconstructionswhere thesubjectfollowstheverb.Forexample, Therearethreestreetsusedbythetraffic. Therearetoomanyusersonthissystem. Thisconstructiontendstoemphasizetheobjectratherthanthesubjectofthesentence.Theflag ‘‘−e’’willcauseSTYLEtoprintallsentencesthatbeginwithanexpletive. 3.DICTION TheprogramDICTIONprintsallsentencesinadocumentcontainingphrasesthatare eitherfrequentlymisusedorindicatewordiness.Theprogram,anextensionofAho’sFGREP[12] stringmatchingprogram,takesasinputafileofphrasesorpatternstobematchedandafileof -7-texttobesearched.Adatabaseofabout450phraseshasbeencompiledasadefaultpatternfile forDICTION.Beforeattemptingtolocatephrases,theprogrammapsuppercaseletterstolower caseandsubstitutesblanksforpunctuation.Sentenceboundariesweredeemedlesscriticalin DICTIONthaninSTYLE,soabbreviationsandotherusesofthecharacter‘‘.’’arenottreated specially.DICTIONbracketsallpatternmatchesinasentencewiththecharacters‘‘[’’‘‘]’’. Althoughmanyofthephrasesinthedefaultdatabasearecorrectinsomecontexts,inothersthey indicatewordiness.Someexamplesofthephrasesandsuggestedalternativesare: PhraseAlternative alargenumberofmany arriveatadecisiondecide collecttogethercollect forthisreasonso pertainingtoabout throughtheuseofbyorwith utilizeuse withtheexceptionofexcept Appendix2containsacompletelistofthedefaultfile.Someoftheentriesareshortformsof problemphrases.Forexample,thephrase‘‘thefact’’isfoundinallofthefollowingandissufficienttopointoutthewordinesstotheuser: PhraseAlternative accountedforbythefactthatcausedby anexampleofthisisthefactthatthus basedonthefactthatbecause despitethefactthatalthough duetothefactthatbecause inlightofthefactthatbecause inviewofthefactthatsince notwithstandingthefactthatalthough EntriesinAppendix2precededby‘‘˜ ’’arenotmatched.SeeSection7fordetailsontheuseof ‘‘˜ ’’. Theusermaysupplyher/hisownpatternfilewiththeflag‘‘−fpatfile’’.Inthiscasethe defaultfilewillbeloadedfirst,followedbytheuserfile.Thismechanismallowsuserstosuppress patternscontainedinthedefaultfileortoincludetheirownpetpeevesthatarenotinthedefault file.Theflag‘‘−n’’willexcludethedefaultfilealtogether.Inconstructingapatternfile,blanks shouldbeusedbeforeandaftereachphrasetoavoidmatchingsubstringsinwords.Forexample, tofindalloccurrencesoftheword‘‘the’’,thepattern‘‘the’’shouldbeused.Theblankscause onlytheword‘‘the’’tobematchedandnotthestring‘‘the’’inwordslikethere,other,andtherefore.Onesideeffectofsurroundingthewordswithblanksisthatwhentwophrasesoccurwithoutinterveningwords,onlythefirstwillbematched. 4.EXPLAIN Thelastprogram,EXPLAIN,isaninteractivethesaurusforphrasesfoundbyDICTION. TheusertypesoneofthephrasesbracketedbyDICTIONandEXPLAINrespondswithsuggested substitutionsforthephrasethatwillimprovethedictionofthedocument. 5.Results 5.1.STYLE Togetbaselinestatisticsandchecktheprogram’saccuracy,weranSTYLEon20technical documents.Therewereatotalof3287sentencesinthesample.Theshortestdocumentwas67 -8-Table1 TextStatisticson20TechnicalDocuments variableminimummaximummeanstandarddeviation ReadabilityKincaid9.516.913.32.2 automated9.017.413.32.5 Cole-Liau10.016.012.71.8 Flesch8.917.014.42.2 sentenceinfo.avsentlength15.530.321.64.0 avwordlength4.615.635.08.29 avnonfunctionlength5.727.306.52.45 shortsent23%46%33%5.9 longsent7%20%14%2.9 sentencetypessimple31%71%49%11.4 complex19%50%33%8.3 compound2%14%7%3.3 compound-complex2%19%10%4.8 verbtypestobe26%64%44.7%10.3 auxiliary10%40%21%8.7 infinitives8%24%15.1%4.8 passives12%50%29%9.3 wordusageprepositions10.1%15.0%12.3%1.6 conjunction1.8%4.8%3.4%.9 adverbs1.2%5.0%3.4%1.0 nouns23.6%31.6%27.8%1.7 adjectives15.4%27.1%21.1%3.4 pronouns1.2%8.4%2.5%1.1 nominalizations2%5%3.3%.8 sentenceopenersprepositions6%19%12%3.4 adverbs0%20%9%4.6 subject56%85%70%8.0 verbs0%4%1%1.0 subordinatingconj1%12%5%2.7 conjunctions0%4%0%1.5 expletives0%6%2%1.7 sentenceslong;thelongest339sentences.Thedocumentscoveredawiderangeofsubjectmatter, includingtheoreticalcomputing,physics,psychology,engineering,andaffirmativeaction.Table1 givestherange,median,andstandarddeviationofthevariousstylemeasures.Asyouwillnote mostofthemeasurementshaveafairlywiderangeofvaluesacrossthesampledocuments. Asacomparison,Table2givesthemedianresultsfortwodifferenttechnicalauthors,asampleofinstructionalmaterial,andasampleoftheFederalistPapers.Thetwoauthorsshowsimilar styles,althoughauthor2usessomewhatshortersentencesandlongerwordsthanauthor1. Author1usesalltypesofsentences,whileauthor2preferssimpleandcomplexsentences,using fewcompoundorcompound-complexsentences.Theothermajordifferenceinthestylesofthese authorsisthelocationofsubordination.Author1seemstopreferembeddedortrailingsubordination,whileauthor2beginsmanysentenceswiththesubordinateclause.Thedocumentstested forbothauthors1and2weretechnicaldocuments,writtenforatechnicalaudience.Theinstructionaldocuments,whicharewrittenforcraftspeople,varysurprisinglylittlefromthetwotechnicalsamples.Thesentencesandwordsarealittlelonger,andtheycontainmanypassiveandauxiliaryverbs,fewadverbs,andalmostnopronouns.Theinstructionaldocumentscontainmany imperativesentences,sotherearemanysentencewithverbopeners.ThesampleofFederalist Paperscontrastswiththeothersamplesinalmosteveryway.

-9-Table2 TextStatisticsonSingleAuthors variableauthor1author2inst.FED readabilityKincaid11.010.310.816.3 automated11.010.311.917.8 Coleman-Liau9.310.110.212.3 Flesch10.310.710.115.0 sentenceinfoavsentlength22.6419.6122.7831.85 avwordlength4.474.664.654.95 avnonfunctionlength5.645.926.046.87 shortsent35%43%35%40% longsent18%15%16%21% sentencetypessimple36%43%40%31% complex34%41%37%34% compound13%7%4%10% compound-complex16%8%14%25% verbtypetobe42%43%45%37% auxiliary17%19%32%32% infinitives17%15%12%21% passives20%19%36%20% wordusageprepositions10.0%10.8%12.3%15.9% conjunctions3.2%2.4%3.9%3.4% adverbs5.05%4.6%3.5%3.7% nouns27.7%26.5%29.1%24.9% adjectives17.0%19.0%15.4%12.4% pronouns5.3%4.3%2.1%6.5% nominalizations1%2%2%3% sentenceopenersprepositions11%14%6%5% adverbs9%9%6%4% subject65%59%54%66% verb3%2%14%2% subordinatingconj8%14%11%3% conjunction1%0%0%3% expletives3%3%0%3% 5.2.DICTION InthefewweeksthatDICTIONhasbeenavailabletousersabout35,000sentenceshave beenrunwithabout5,000stringmatches.Theauthorsusingtheprogramseemtomakethesuggestedchangesabout50-75%ofthetime.Todate,almost200ofthe450stringsinthedefault filehavebeenmatched.Althoughmostofthesephrasesarevalidandcorrectinsomecontexts, the50-75%changerateseemstoshowthatthephrasesareusedmuchmoreoftenthanconcise dictionwarrants. 6.Accuracy 6.1.SentenceIdentification ThecorrectnessoftheSTYLEoutputonthe20documentsamplewascheckedindetail. STYLEmisidentified129sentencefragmentsassentencesandincorrectlyjoinedtwoormoresentences75timesinthe3287sentencesample.Theproblemswereusuallybecauseofnonstandard formattingcommands,unknownabbreviations,orlistsofnon-sentences.Animpossiblylongsen-tencefoundasthelongestsentenceinthedocumentusuallyistheresultofalonglistofnon-sentences.

-10-6.2.SentenceTypes Stylecorrectlyidentifiedsentencetypeon86.5%ofthesentencesinthesample.Thetype distributionofthesentenceswas52.5%simple,29.9%complex,8.5%compoundand9%compound-complex.Theprogramreported49.5%simple,31.9%complex,8%compoundand10.4% compound-complex.Lookingattheerrorsontheindividualdocuments,thenumberofsimple sentenceswasunder-reportedbyabout4%andthecomplexandcompound-complexwereover-reportedby3%and2%,respectively.Thefollowingmatrixshowstheprogramsoutputvs.the actualsentencetype. ProgramResults simplecomplexcompoundcomp-complex Actualsimple15661324917 Sentencecomplex47892665 Typecompound40620723 comp-complex0525249 Thesystem’sinabilitytofindimperativesentencesseemstohavelittleeffectonmostofthe stylestatistics.Adocumentwithhalfofitssentencesimperativewasrun,withandwithoutthe imperativeendmarker.Theresultswereidenticalexceptfortheexpectederrorsofnotfinding verbsassentenceopeners,notcountingtheimperativesentences,andaslightdifference(1%)in thenumberofnounsandadjectivesreported. 6.3.WordUsage TheaccuracyofidentifyingwordtypesreflectsthatofPARTS,whichisabout95%correct. Thelargestsourceofconfusionisbetweennounsandadjectives.Theverbcountswerecheckedon about20sentencesfromeachdocumentandfoundtobeabout98%correct. 7.TechnicalDetails 7.1.FindingSentences Theformattingcommandsembeddedinthetextincreasethedifficultyoffindingsentences. Notalltextinadocumentisinsentenceform;thereareheadings,tables,equationsandlists,for example.Headingslike‘‘FindingSentences’’aboveshouldbediscarded,notattachedtothenext sentence.However,sincemanyofthedocumentsareformattedtobephototypeset,andcontain fontchanges,whichusuallyoperateonthemostimportantwordsinthedocument,discardingall formattingcommandsisnotcorrect.Toimprovetheprograms’abilitytofindsentenceboundaries,thedeformattingprogram,DEROFF[13],hasbeengivensomeknowledgeoftheformatting packagesusedontheUNIXoperatingsystem.DEROFFwillnowdothefollowing: 1.Suppressallformattingmacrosthatareusedfortitles,headings,author’sname,etc. 2.Suppresstheargumentstothemacrosfortitles,headings,author’sname,etc. 3.Suppressdisplays,tables,footnotesandtextthatiscenteredorinno-fillmode. 4.Substituteaplaceholderforequationsandcheckforhiddenendmarkers.Theplaceholder isnecessarybecausemanytypistsandauthorsusetheequationsettertochangefontson importantwords.Forthisreason,headerfilescontainingthedefinitionoftheEQNdelimitersmustalsobeincludedasinputtoSTYLE.Endmarkersareoftenhiddenwhenan equationendsasentenceandtheperiodistypedinsidetheEQNdelimiters. 5.Adda"."afterlists.Iftheflag−mlisalsoused,alllistsaresuppressed.Thisisaseparate flagbecauseofthevarietyofwaysthelistmacrosareused.Often,listsaresentencesthat shouldbeincludedintheanalysis.Theusermustdeterminehowlistsareusedinthedocumenttobeanalyzed. BothSTYLEandDICTIONcallDEROFFbeforetheylookatthetext.Theusershould supplythe−mlflagifthedocumentcontainsmanylistsofnon-sentencesthatshouldbeskipped.

-11-7.2.DetailsofDICTION TheprogramDICTIONisbasedonthestringmatchingprogramFGREP.FGREPtakesas inputafileofpatternstobematchedandafiletobesearchedandoutputseachlinethatcontains anyofthepatternswithnoindicationofwhichpatternwasmatched.Thefollowingchangeshave beenaddedtoFGREP: 1.ThebasicunitthatDICTIONoperatesonisasentenceratherthanaline.Eachsentence thatcontainsoneofthepatternsisoutput. 2.Uppercaselettersaremappedtolowercase. 3.Punctuationisreplacedbyblanks. 4Allpatternmatchesinthesentencearefoundandsurroundedwith‘‘[’’‘‘]’’. 5.Amethodforsuppressingastringmatchhasbeenadded.Anypatternthatbeginswith‘‘˜ ’’ willnotbematched.Becausethematchingalgorithmfindsthelongestsubstring,thesuppressionofamatchallowswordsinsomecorrectcontextsnottobematchedwhileallowing thewordinanothercontexttobefound.Forexample,theword‘‘which’’isoftenincorrectly usedinsteadof‘‘that’’inrestrictiveclauses.However,‘‘which’’isusuallycorrectwhenprecededbyaprepositionor‘‘,’’.Thedefaultpatternfilesuppressesthematchofthecommon prepositionsoradoubleblankfollowedby‘‘which’’andthereforematchesonlythesuspect uses.Thedoubleblankaccountsforthereplacedcomma. 8.Conclusions Asystemofwritingtoolsthatmeasuresomeoftheobjectivecharacteristicsofwritingstyle hasbeendeveloped.Thetoolsaresufficientlygeneralthattheymaybeappliedtodocumentson anysubjectwithequalaccuracy.Althoughthemeasurementsareonlyofthesurfacestructureof thetext,theydopointoutproblemareas.Inadditiontohelpingwritersproducebetterdocuments,theseprogramsmaybeusefulforstudyingthewritingprocessandfindingotherformulae formeasuringreadability.

-12-References 1.L.L.Cherry,‘‘PARTS-ASystemforAssigningWordClassestoEnglishText,’’submitted CommunicationsoftheACM. 2.B.W.KernighanandJ.R.Mashey,‘‘TheUNIXProgrammingEnvironment,’’Software− Practice&Experience,9,1-15(1979). 3.G.R.Klare,‘‘AssessingReadability,’’ReadingResearchQuarterly,1974-1975,10,62-102. 4.E.A.SmithandP.Kincaid,‘‘Derivationandvalidationoftheautomatedreadabilityindex forusewithtechnicalmaterials,’’HumanFactors,1970,12,457-464. 5.J.P.Kincaid,R.P.Fishburne,R.L.Rogers,andB.S.Chissom,‘‘Derivationofnewreadabilityformulas(AutomatedReadabilityIndex,Fogcount,andFleschReadingEaseFormula)forNavyenlistedpersonnel,’’NavyTrainingCommandResearchBranchReport8-75, Feb.,1975. 6.M.ColemanandT.L.Liau,‘‘AComputerReadabilityFormulaDesignedforMachineScoring,’’JournalofAppliedPsychology,1975,60,283-284. 7.R.Flesch,‘‘ANewReadabilityYardstick,’’JournalofAppliedPsychology,1948,32,221-233. 8.E.U.Coke,privatecommunication. 9.D.W.Ewing,WritingforResults,JohnWiley&Sons,Inc.,NewYork,N.Y.(1974). 10.G.Leggett,C.D.MeadandW.Charvat,Prentice-HallHandbookforWriters,SeventhEdi-tion,Prentice-HallInc.,EnglewoodCliffs,N.J.(1978). 11.E.B.Coleman,‘‘LearningofProseWritteninFourGrammaticalTransformations,’’Journal ofAppliedPsychology,1965,vol.49,no.5,pp.332-341. 12A.V.AhoandM.J.Corasick,‘‘EfficientStringMatching:anaidtoBibliographicSearch,’’ CommunicationsoftheACM,18,(6),333-340,June1975. 13.BellLaboratories,‘‘UNIXTIME-SHARINGSYSTEM:UNIXPROGRAMMER’SMAN-UAL,’’SeventhEdition,Vol.1(January1979).

-13-Appendix1 STYLEAbbreviations a.d. A.M. a.m. b.c. Ch. ch. ckts. dB. Dept. dept. Depts. depts. Dr. Drs. e.g. Eq. eq. etal. etc. Fig. fig. Figs. figs. ft. i.e. in. Inc. Jr. jr. mi. Mr. Mrs. Ms. No. no. Nos. nos. P.M. p.m. Ph.D. Ph.d. Ref. ref. Refs. refs. St. vs. yr.

-14-Appendix2 DefaultDICTIONPatterns agreatdealof alargenumberof alotof amajorityof aneedfor anumberof aparticularpreferencefor apreferencefor asmallnumberof atendencyto abovementioned absolutelycomplete absolutelyessential accomplished accordingly activate actual addedincrements adequateenough advent affordanopportunity aggregate allof allthroughout alongtheline anindicationof analyzation andetc andor anotheradditional anyandall arriveata asamatteroffact asamethodof asgoodorbetterthan asofnow asper asregards asrelatedto asto assistance assistanceto assistanceto assumingthat atalaterdate atabout atabove atalltimes atanearlydate atbelow atthepresent atthetimewhen atthispointintime atthistime atwhichtime atyourearliestconvenience authorization awful basicfundamentals basically becognizantof beingas beingthat briefinduration bringtoaconclusion butthat butwhat bymeansof bytheuseof carryoutexperiments centerabout centeraround centerportion checkinto checkon checkupon circlearound closeproximity collaboratetogether collecttogether combinetogether cometoanend commence commonaccord compensation completelyeliminated comprise concerning conductaninvestigationof conjecture connectup consensusofopinion consequentresult consolidatetogether construct contemplate continueon continuetoremain couldof countup coupletogether debateabout decideon deleteriouseffect demean demonstrate depreciateinvalue deservingof desirablebenefits desirousof differentthan discontinue disutility divideup doubtbut dueto dulynoted duringthetimethat eachandevery earlybeginnings effectuate emotionalfeelings emptyout enclosedherein enclosedherewith endresult endup endeavor enterin enterinto enthused entirelycomplete equallygoodas essentially eventuate everynowandthen exactlyidentical experiencingdifficulty fabricate faceupto facilitate factsandfigures fastinaction fearfulof fearfulthat fewinnumber fileaway finalcompletion finalending finaloutcome finalresult finalize finditinterestingtoknow firstandforemost firstbeginnings firstinitiated firstly followafter followingafter forthepurposeof forthereasonthat forthesimplereasonthat forthisreason foryourinformation fromthepointofviewof fullandcomplete generallyagreed goodand gotto gratuitous greatlyminimize headup helpbut helpsintheproductionof hopeful ifandwhen ifatallpossible impact implement importantessentials importantly inalargemeasure inapositionto inaccordance inadvanceof inagreementwith inallcases inbackof inbehalfof inbehind inbetween incase incloseproximity inconflictwith inconjunctionwith inconnectionwith infact inlargemeasure inmanycases inmostcases inmyopinionIthink inorderto inrarecases inreferenceto inregardto inregardsto inrelationwith inshortsupply insize intermsof intheamountof inthecaseof inthecourseof intheevent inthefieldof intheformof intheinstanceof intheinterim inthelastanalysis inthematterof inthenearfuture intheneighborhoodof inthenottoodistantfuture intheproximityof intherangeof inthesamewayasdescribed intheshapeof inthevicinityof inthiscase inviewofthe inviolationof inasmuchas indicate indicativeof initialize initiate injuriousto inquire insideof institutea intentsandpurposes intermingle irregardless isdefinedas isusedtocontrol iswhen iswhere itisincumbent itstandstoreason itwasnotedthatif jointcooperation jointpartnership justexactly kindof knowabout lastbutnotleast lateron leavingoutofconsideration liable linkup literally littledoubtthat loseouton lotsof mainessentials makea makeadjustmentsto makean makeapplicationto makecontactwith makementionof makeoutalistof maketheacquaintanceof maketheadjustment manner maximumpossible meaningful meetupwith meltdown meltup methodology mightof minimizeasfaraspossible minorimportance missouton modification -15-morepreferable mostunique mustof mutualcooperation necessaryrequisite necessitate needfor nice notbeun notinapositionto notofahighorderofaccuracy notun notwithstanding ofconsiderablemagnitude ofthat oftheopinionthat offof onafewoccasions onaccountof onbehalfof onthegroundsthat ontheoccasion onthepartof oneofthe openup operatestocorrect outsideof overwith overall pasthistory perceptiveof performameasurement performthemeasurement permitsthereductionof personalize pertainingto physicalsize planahead planforthefuture planinadvance planon presentaconclusion presentareport presently priorto prioritize proceedto procure productiveof prolongtheduration protrudeoutfrom providedthat pursuantto puttousein rangeallthewayfrom reasonisbecause reasonwhy recuragain reducedown referback referencetothis reflectiveof regarding regretful reinitiate relativeto repeatagain representativeof resultanteffect resumeagain retreatback returnagain returnback revertback sealoff seemsapparent sendacommunication shortspaceoftime shouldof singleunit situation soasto sortof spellout stillcontinue stillremain subsequent substantiallyinagreement succeedin suggestiveof superiorthan surroundingcircumstances takeappropriate takecognizanceof takeintoconsideration termedas terminate termination theauthor theauthors thecasethat thefact theforegoing theforeseeablefuture thefullestpossibleextent themajorityof thenature thenecessityof theonlydifferencebeingthat theorderof thepointthat thetruthis therearenotmany throughthemediumof throughtheuseof throughouttheentire timeinterval tosummarizetheabove totaleffectofallthis totality transpire truefacts tryand ultimateend underaseparatecover underdateof underseparatecover underthenecessityto underlyingpurpose undertakeastudy uniformlyconsistent unique untilsuchtimeas uptothistime upshot utilize very verycomplete veryunique vital which withaviewto withreferenceto withregardto withtheexceptionof withtheobjectof withtheresultthat withthisinmind,itisclearthat withintherealmofpossibility withoutfurtherdelay worthwhile wouldof ingbehavior wise ˜ which ˜ aboutwhich ˜ afterwhich ˜ atwhich ˜ betweenwhich ˜ bywhich ˜ forwhich ˜ fromwhich ˜ inwhich ˜ intowhich ˜ ofwhich ˜ onwhich ˜ onwhich ˜ overwhich ˜ throughwhich ˜ towhich ˜ underwhich ˜ uponwhich ˜ withwhich ˜ withoutwhich ˜ clockwise ˜ likewise ˜ otherwise